OpenStudy (anonymous):

Find the coefficient x^7 in \[\huge (ax^{2}+\frac{ 1 }{ bx })^{11}\] and of x^ -7 in \[\huge (ax-\frac{ 1 }{ bx^{2} })^{11}\] and find the relation between "a" and "b" so that these coefficients are equal (where \[a,b \neq 0\]

Parth (parthkohli):

OK - I don't know what you did, and I might have taken the longer route, but what is the answer given?

OpenStudy (anonymous):

ab=1 or a=1/b

OpenStudy (anonymous):

in the expansion of \[\left( x+a \right)^n\] General term is \[T _{r+1} =C_{r}^{n}x ^{n-r}a ^{r}\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[here x=ax^2,a= \frac{ 1 }{ bx }\]

OpenStudy (anonymous):

yes agreed

OpenStudy (anonymous):

\[T _{r+1}=C _{r}^{11}\left( a x^2 \right)^{11-r}\left( \frac{ 1 }{ bx } \right)^r\] \[=C _{r}^{11}a ^{11-r}x ^{22-2r}\frac{ 1 }{ b^rx^r }=C _{r}^{11}a ^{11-r}b ^{-r}x ^{22-2r-r}\] \[=C _{r}^{11}a ^{11-r}b ^{-r}x ^{22-3r}\] put 22-3r=7 3r=22-7=15 r=5 now find the coefficient.

OpenStudy (anonymous):

i made a mistake i didn't take x^2

OpenStudy (anonymous):

It is \[C _{5}^{11}a ^{11-5}b ^{-5}=?\]

OpenStudy (anonymous):

\[\huge C_{5}^{11}\]

OpenStudy (anonymous):

that's the coefficient

OpenStudy (anonymous):

no, a and b are also constants,any thing other than x and its power is coefficient.

OpenStudy (anonymous):

\[\huge C _{5}^{11}a^{6}b^{-5}\]

OpenStudy (anonymous):

well negative 1/bx isn't there

OpenStudy (anonymous):

its postive

OpenStudy (anonymous):

now i have seen it is in second question.

OpenStudy (anonymous):

in second \[a=-\frac{ 1 }{ bx^2 }\]

OpenStudy (anonymous):

we are doing first right

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

okay i understand the situation , well , then i get \[\huge C _{5}^{11}a^{6}b^{-5}\]

OpenStudy (anonymous):

correct.

OpenStudy (anonymous):

\[C _{5}^{11}=\frac{ 11*10*9*8*7 }{ 5*4*3*2*1 }=?\]

OpenStudy (anonymous):

132

OpenStudy (anonymous):

now solve second as we solved first then equate coefficients.

OpenStudy (anonymous):

doing it

OpenStudy (anonymous):

I think 132 is not correct.

OpenStudy (anonymous):

\[\left(a x^2+\frac{1}{b x}\right)^{11}=a^{11} x^{22}+\frac{11 a^{10} x^{19}}{b}+\frac{55 a^9 x^{16}}{b^2}+\frac{165 a^8 x^{13}}{b^3}+\frac{330 a^7 x^{10}}{b^4}+\\\frac{462 a^6 x^7}{b^5}+\frac{462 a^5 x^4}{b^6}+\frac{330 a^4 x}{b^7}+\frac{165 a^3}{b^8 x^2}+\frac{55 a^2}{b^9 x^5}+\frac{11 a}{b^{10} x^8}+\frac{1}{b^{11} x^{11}}\]

OpenStudy (anonymous):

it is 462 so coefficient is \[462~a^6~b ^{-5}\]

OpenStudy (anonymous):

for second x=ax

OpenStudy (anonymous):

\[\left(a x-\frac{1}{b x^2}\right)^{11}=a^{11} x^{11}-\frac{11 a^{10} x^8}{b}+\frac{55 a^9 x^5}{b^2}-\frac{165 a^8 x^2}{b^3}+\frac{330 a^7}{b^4 x}-\frac{462 a^6}{b^5 x^4}+\\\frac{462 a^5}{b^6 x^7}-\frac{330 a^4}{b^7 x^{10}}+\frac{165 a^3}{b^8 x^{13}}-\frac{55 a^2}{b^9 x^{16}}+\frac{11 a}{b^{10} x^{19}}-\frac{1}{b^{11} x^{22}} \]

OpenStudy (anonymous):

\[ \frac{462 a^5}{b^6}=\frac{462 a^6}{b^5} \] you get a b=1

OpenStudy (anonymous):

that is a long procedure ,better take a general term as we did in first.

OpenStudy (anonymous):

Sorry I miss the minus sign we need ab=-1

OpenStudy (anonymous):

And the equation to be solved \[ \frac{462 a^5}{b^6}=-\frac{462 a^6}{b^5} \]

OpenStudy (anonymous):

The second equation becomes the first, if we change a=-1/b and b =-1/a meaning ab=-1

OpenStudy (skullpatrol):

@eliassaab do you have any advice for keeping track of minus signs?

OpenStudy (anonymous):

It alternates form + to -1 if you apply the binomial theorem to the power of the difference.

OpenStudy (skullpatrol):

I meant while working on math questions in general @eliassaab ?

OpenStudy (anonymous):

You cannot make a statement about some general question like this one

OpenStudy (skullpatrol):

Thank you